# One Model to Rank Them All: Unifying Online Advertising with End-to-End Learning

### v1

@@ Abstract.
Modern industrial advertising systems commonly employ Multi-stage Cascading Architectures (MCA) to balance computational efficiency with ranking accuracy. However, this approach presents two fundamental challenges: (1) performance inconsistencies arising from divergent optimization targets and capability differences between stages, and (2) failure to account for advertisement externalities - the complex interactions between candidate ads during ranking. These limitations ultimately compromise system effectiveness and reduce platform profitability. In this paper, we present UniROM, an end-to-end generative architecture that Unifies online advertising Ranking as One Model. UniROM replaces cascaded stages with a single model to directly generate optimal ad sequences from the full candidate ad corpus in location-based services (LBS). The primary challenges associated with this approach stem from high costs of feature processing and computational bottlenecks in modeling externalities of large-scale candidate pools. To address these challenges, UniROM introduces an algorithm and engine co-designed hybrid feature service to decouple user and ad feature processing, reducing latency while preserving expressiveness. To efficiently extract intra- and cross-sequence mutual information, we propose RecFormer with an innovative cluster-attention mechanism as its core architectural component. Furthermore, we propose a bi-stage training strategy that integrates pre-training with reinforcement learning-based post-training to meet sophisticated platform and advertising objectives. Extensive offline evaluations on public benchmarks and large-scale online A/B testing on industrial advertising platform have demonstrated the superior performance of UniROM over state-of-the-art MCAs.
@@ 1. INTRODUCTION
In this paper, we present UniROM, a novel framework that Unifies online advertising Ranking as One Model in location-based services (LBS). By narrowing the candidate set to ads within the same city (reducing its size to $\sim 10^{5}$), UniROM can perform fine-grained contextual modeling across the entire candidate space, enabling both more accurate item representations and sufficient exploration of user interests. Furthermore, this architecture conceptualizes online advertising systems as a unified generative process, thereby eliminating the inherent goal conflicts between different pipeline stages. Nevertheless, the proposed architecture is confronted with three fundamental challenges:
@@
To minimize feature storage and transmission overhead while maintaining expressiveness, we develop a novel Hybrid Feature Service (HFS) that optimizes computational efficiency by decoupling user and ad features. Additionally, by enabling batch processing of multiple candidates in a single pass, HFS significantly amortizes computational and I/O costs, thereby improving system scalability. Furthermore, we present RecFormer, an transformative recommendation framework that leverages an innovative cluster-attention mechanism to simultaneously model users’ deep interests and contextual externalities while maintaining computational efficiency. To substantially enhance inference speed without compromising key platform optimization objectives, such as user engagement and revenue performance, UniROM incorporates an AucFormer module that generates advertising sequences through non-autoregressive processing. Finally, we explore a multi-stage training strategy to ensure consistency with platform objectives and auction constraints. To summarize, our main contributions are in three-fold:
@@
Figure 2. Architecture Overview of UniROM, showcasing its key components including the Hybrid Feature Service, RecFormer and AucFormer.
@@ 3. METHODOLOGY
As illustrated in Figure 2, UniROM operates through three integrated components: 1) The Hybrid Feature Service (HFS) efficiently extracting and processing large-scale, fine-grained features; 2) RecFormer modeling deep user interests and candidate externalities across the full candidate set; and 3) AucFormer optimizing ranking to align with advertising platform objectives. Finally, we employ a bi-stage training strategy, wherein the pre-training phase focuses on modeling user preferences, while the post-training phase optimizes for profitability.
@@ 3.4. AucFormer
Traditional auction mechanisms like GSP (Edelman et al., 2007) fail to model externalities effectively and lack deep integration with machine learning technologies. In this paper, UniROM introduces the AucFormer framework, comprising 1) a non-autoregressive (NAR) generator that employs a matching mechanism to simultaneously predict allocation probabilities for candidate ads across multiple interface slots [25], and 2) a permutation-aware evaluator that leverages exposure order information to more accurately estimate user engagement metrics, e.g. pCTR, pCVR.
@@ 4. TRAINING AND OPTIMIZATION
In this section, we present a two-stage training framework inspired by large language model optimization (Ouyang et al., 2022). The first stage, pre-training, focuses on aligning UniROM with user engagement signals such as clicks and purchases. Subsequently, the post-training stage employs reinforcement learning to optimize UniROM’s response to auction feedback and guide the payment network under predefined economic constraints.
@@ 5. EXPERIMENTS
In this section, we evaluate the effectiveness of UniROM using offline experiments and online A/B testing.
@@ 5.1.1. Dataset
We provide empirical evidence for the effectiveness of UniROM on the industrial dataset Meituan  [^4] Due to the need to protect business secrets, some transformations were applied to the results. These transformations were carefully designed to maintain the statistical properties while ensuring that no sensitive business - related information could be reverse - engineered from the transformed results. from a well-known local life platform that leverages Location-Based Services (LBS) to provide users with convenient and tailored services. The industrial dataset is constructed from the real logs of advertising platform from April 2024 to October 2024. It contains 200 million requests from more than 2 million users across nearly 10 million ads. The first 200 days are used for pre-training, a randomly sampled 50 days subset is employed fro post-training, and the last 14 days are used as the test set. In the LBS scenario, the maximum number of valid candidate ads in each sample is $N=100,000$. To enhance training efficiency, popularity sampling is utilized during training, with $N_{s}=2995$ and $K=5$. During testing, inference is conducted using all candidate ads for each sample, where $N=100,000$ and $K=5$ are employed.
@@ 5.1.3. Baselines
We compare UniROM with the following two representative architectures, which are widely used in industry: 1) MCA. The Multi-stage Cascading Architecture is a common design for online advertising systems. To implement this architecture effectively, we employ four representation methods: SASREC (Kang and McAuley, 2018) for recall, DSSM (Huang et al., 2013) for pre-ranking, DIN (Zhou et al., 2018) for ranking, and GSP (Edelman et al., 2007) for auction. 2) FS-LTR. Full Stage Learning to Rank (Zheng et al., 2024b) is a unified training framework designed for multi-stage recommendation systems. It leverages relabeled data from all stages to train models in MCA, ensuring that the top-ranked items are more likely to pass through subsequent stages and align with user interests. To ensure effectiveness and comparability of FS-LTR, we use the same architecture as MCA, incorporating FS-LTR’s sample re-labeling technique during training to enhance consistency.
@@ 5.1.4. Hyperparameters
For MCA and FS-LTR, we follows the hyperparameters setting in (Liu et al., 2024). For UniROM, we tried different hyperparameters using grid search. Due to space limitations, only the most optimal parameters are presented in this paper. The clustering approach $G$ employs an adaptive clustering mechanism, the hidden layers of the MLP are 128 and 32, the learning rate is $10^{-3}$, the batch size is $128$ and the optimizer is Adam. the user behavior sequence length $L=1000$, the dimension size $d=128$, the layer number of GCF and MIF $m=6$, the layer number of RecFormer $m_{e}=3$, the number of interval layers in MIF $m_{k}=2$, the number of attention heads $N_{h}=4$ and the number of adaptive clusters $N_{c}=128$.
@@ 5.2. Offline Experiments
Table 1 summarizes the results on industrial dataset from the offline experiments. Each offline experiment is repeated $5$ times with different random seeds and each result is presented in the form of mean ± standard. The experimental results yield the following observations. On the industrial dataset, UniROM improves over the state-of-the-art baselines in AUC, Recall@50, eCTR, eRPM, and $\Psi$ respectively, demonstrating the superiority of UniROM. Specifically, UniROM achieves a Recall@50 of 0.513, representing a significant lift of 20.4% [^5] Lift percentage means the improvement of UniROM over the best baselines over the strongest baseline FS-LTR. The AUC increases to 0.754 (+1.48%), and eCTR improves +8.3%, demonstrating superior user modeling and ranking capabilities. In terms of platform revenue, UniROM achieves an eRPM of 217.1, which is 11.4% higher than FS-LTR. Most notably, the IC metric $\Psi$ is dramatically reduced to 2.3%, compared to 9.1% and 9.4% for FS-LTR and MCA respectively, indicating much stronger incentive compatibility and robustness to auction constraints.
@@ 5.3. Ablation Study
To verify the effectiveness of UniROM’s various design considerations, we construct three variants:
  • $\textbf{UniROM}_{-{gcf}}$ removes the GCF and uses point-wise pCTR instead of set-aware pCTR for next tasks.
  • $\textbf{UniROM}_{-{mif}}$ removes the MIF and only uses DIN to capture user’s interests from user’s historical behavior sequence.
  • $\textbf{UniROM}_{-auf}$ removes the AucFormer and uses GSP mechanism for completing ad allocation and payment.
Judging from the online experimental results in Table 2, we have the following findings: 1) The variant without GCF performs worse than UniROM. This phenomenon proves that our proposed GCF can effectively model the global set-aware externalities and help UniROM to achieve better performance. Especially in our LBS scenarios, the benefits of this global externalities modeling are further enhanced. 2) The experimental results of $\textbf{UniROM}_{-{mif}}$ are worse than UniROM, this supports that the deep modeling of user behavior sequences is crucial for both prediction and generation tasks. (3) The performance gap between w/ and w/o AucFormer is obvious. This indicates that AucFormer structure and RLAF training approach can effectively align with auction preferences, resulting in improved ad sequence generation and overall performance.
@@ 5.4. Cross Features Analysis
UniROM 0.754 0.005
UniROM w/o CF 0.751 (-0.3%) 0.011 (+0.6%)
UniROM-${}_{\text{-CA}}$ w/o CF 0.745 (-0.9%) 0.036 (+3.1%)
UniROM-${}_{\text{-TA}}$ w/o CF 0.743(-1.1%) 0.047 (+4.2%)
UniROM-${}_{\text{-mif}}$ w/o CF 0.737 (-1.7%) 0.048 (+4.4%)
@@ 5.5. Scaling Law
In UniROM, each module uses stacked blocks based on cluster-attention as core structure to ensure scalability. Here, we vary the number of blocks across modules to study scaling laws. Besides, we Results in Figure 5 reveal:
@@
2) The figure on the right reveals a clear positive correlation between computational cost and model performance in UniROM. The upward-trending linear fit (brown dashed line) indicates that increased computational resources consistently enhance recommendation accuracy, though the marginal gains gradually diminish at higher FLOPs levels. This suggests that while scaling computation remains an effective strategy for performance improvement within the tested range, the model’s architecture or data constraints may eventually limit further gains—mirroring the saturation pattern observed in block stacking. The relationship underscores the importance of balancing computational investment against diminishing returns in practical deployment scenarios.
@@ 5.6. Online Results
To verify UniROM’s effectiveness in the real-world, we compare UniROM with the fully deployed MCA in industrial advertising system through online A/B tests. Table 4 presents the results of online A/B testing conducted from November 18 to November 24, 2024. Experimental results demonstrate 3.8% improvements in CTR, 11.2% in RPM, and 3.1% in ROI respectively, indicating UniROM’s effectiveness in enhancing user experience, platform revenue, and advertiser utility. Notably, the harmonized gains in user engagement and advertiser’s ROI suggest that revenue growth stems not from payment inflation, but from enhanced modeling capabilities.
We also analyzed the online inference computational complexity of UniROM and MCA. UniROM processes hundreds of times more candidate ads than MCA’s ranking model with only a 5 ms average increase (2.2% relatively) in online response time (RT) per request). The complexity is discussed in Appendix A.1.
@@ 6. CONCLUSIONS
In this paper, we propose UniROM that Unifies the entire ad ranking and allocation pipeline as One Model. Our key innovations include: 1) a hybrid feature service that decouples user and ad feature processing to reduce latency while preserving expressiveness; 2) an transformative recommendation framework that leverages an innovative cluster-attention mechanism to efficiently model users’ deep interests and contextual externalities; and 3) a bi-stage training strategy align with user preference and platform objectives. Extensive experiments on public and industrial datasets demonstrate UniROM’s superiority over state-of-the-art approaches. In the future, we will further explore scaling laws and different training paradigms to enhance UniROM’s capabilities.
@@ A.1. Complexity Analysis
We provide a detailed comparison of computational complexity between MCA and UniROM. All floating-point operations (FLOPs) are calculated based on standard Transformer block operations (Vaswani et al., 2017). Given sequence length $L$, hidden dimension $d$, the total FLOPs of a standard Transformer block consist of three components:
@@
UniROM Complexity. The FLOPs of UniROM contains three modules:
@@
Finally, the total Flops of UniROM is calcalated as:
(33) $\text{FLOPs}_{{UniROM}}=$ $\text{FLOPs}_{{gcf}}+\text{FLOPs}_{{mif}}+\text{FLOPs}_{{auf}}.$
Both the computational complexities of MCA and UniROM are proportional to the number of candidates($N$). With $N_{r}=\alpha N$, $m_{r}=m$, and $m=2m_{c}$, when $N$ is large, the ratio of proportions can be approximated as follows:
$\frac{\text{FLOPs}_{{UniROM}}}{\text{FLOPs}_{{mca}}}\approx\frac{2m_{k}NLd^{2}}{m_{r}N_{r}4L^{2}d}$
In our case, with $\alpha=0.033$, this results in a ratio of $0.97$. The slight difference in proportions leads to a significant FLOPs gap as the number of candidates increases, demonstrating UniROM’s capability and efficiency within our design.

### v2

# EGA-V1: Unifying Online Advertising with End-to-End Learning

@@ Abstract.
11footnotetext: Equal contribution.22footnotetext: Corresponding authors.
Modern industrial advertising systems commonly employ Multi-stage Cascading Architectures (MCA) to balance computational efficiency with ranking accuracy. However, this approach presents two fundamental challenges: (1) performance inconsistencies arising from divergent optimization targets and capability differences between stages, and (2) failure to account for advertisement externalities - the complex interactions between candidate ads during ranking. These limitations ultimately compromise system effectiveness and reduce platform profitability. In this paper, we present EGA-V1, an end-to-end generative architecture that unifies online advertising ranking as one model. EGA-V1 replaces cascaded stages with a single model to directly generate optimal ad sequences from the full candidate ad corpus in location-based services (LBS). The primary challenges associated with this approach stem from high costs of feature processing and computational bottlenecks in modeling externalities of large-scale candidate pools. To address these challenges, EGA-V1 introduces an algorithm and engine co-designed hybrid feature service to decouple user and ad feature processing, reducing latency while preserving expressiveness. To efficiently extract intra- and cross-sequence mutual information, we propose RecFormer with an innovative cluster-attention mechanism as its core architectural component. Furthermore, we propose a bi-stage training strategy that integrates pre-training with reinforcement learning-based post-training to meet sophisticated platform and advertising objectives. Extensive offline evaluations on public benchmarks and large-scale online A/B testing on industrial advertising platform have demonstrated the superior performance of EGA-V1 over state-of-the-art MCAs.
@@ 1. INTRODUCTION
In this paper, we present EGA-V1, a novel framework that unifies online advertising ranking as one model in location-based services (LBS). By narrowing the candidate set to ads within the same city (reducing its size to $\sim 10^{5}$), EGA-V1 can perform fine-grained contextual modeling across the entire candidate space, enabling both more accurate item representations and sufficient exploration of user interests. Furthermore, this architecture conceptualizes online advertising systems as a unified generative process, thereby eliminating the inherent goal conflicts between different pipeline stages. Nevertheless, the proposed architecture is confronted with three fundamental challenges:
@@
To minimize feature storage and transmission overhead while maintaining expressiveness, we develop a novel Hybrid Feature Service (HFS) that optimizes computational efficiency by decoupling user and ad features. Additionally, by enabling batch processing of multiple candidates in a single pass, HFS significantly amortizes computational and I/O costs, thereby improving system scalability. Furthermore, we present RecFormer, an transformative recommendation framework that leverages an innovative cluster-attention mechanism to simultaneously model users’ deep interests and contextual externalities while maintaining computational efficiency. To substantially enhance inference speed without compromising key platform optimization objectives, such as user engagement and revenue performance, EGA-V1 incorporates an AucFormer module that generates advertising sequences through non-autoregressive processing. Finally, we explore a multi-stage training strategy to ensure consistency with platform objectives and auction constraints. To summarize, our main contributions are in three-fold:
@@
Figure 2. Architecture Overview of EGA-V1, showcasing its key components including the Hybrid Feature Service, RecFormer and AucFormer.
@@ 3. METHODOLOGY
As illustrated in Figure 2, EGA-V1 operates through three integrated components: 1) The Hybrid Feature Service (HFS) efficiently extracting and processing large-scale, fine-grained features; 2) RecFormer modeling deep user interests and candidate externalities across the full candidate set; and 3) AucFormer optimizing ranking to align with advertising platform objectives. Finally, we employ a bi-stage training strategy, wherein the pre-training phase focuses on modeling user preferences, while the post-training phase optimizes for profitability.
@@ 3.4. AucFormer
Traditional auction mechanisms like GSP (Edelman et al., 2007) fail to model externalities effectively and lack deep integration with machine learning technologies. In this paper, EGA-V1 introduces the AucFormer framework, comprising 1) a non-autoregressive (NAR) generator that employs a matching mechanism to simultaneously predict allocation probabilities for candidate ads across multiple interface slots [25], and 2) a permutation-aware evaluator that leverages exposure order information to more accurately estimate user engagement metrics, e.g. pCTR, pCVR.
@@ 4. TRAINING AND OPTIMIZATION
In this section, we present a two-stage training framework inspired by large language model optimization (Ouyang et al., 2022). The first stage, pre-training, focuses on aligning EGA-V1 with user engagement signals such as clicks and purchases. Subsequently, the post-training stage employs reinforcement learning to optimize EGA-V1’s response to auction feedback and guide the payment network under predefined economic constraints.
@@ 5. EXPERIMENTS
In this section, we evaluate the effectiveness of EGA-V1 using offline experiments and online A/B testing.
@@ 5.1.1. Dataset
We provide empirical evidence for the effectiveness of EGA-V1 on the industrial dataset Meituan  [^4] Due to the need to protect business secrets, some transformations were applied to the results. These transformations were carefully designed to maintain the statistical properties while ensuring that no sensitive business - related information could be reverse - engineered from the transformed results. from a well-known local life platform that leverages Location-Based Services (LBS) to provide users with convenient and tailored services. The industrial dataset is constructed from the real logs of advertising platform from April 2024 to October 2024. It contains 200 million requests from more than 2 million users across nearly 10 million ads. The first 200 days are used for pre-training, a randomly sampled 50 days subset is employed fro post-training, and the last 14 days are used as the test set. In the LBS scenario, the maximum number of valid candidate ads in each sample is $N=100,000$. To enhance training efficiency, popularity sampling is utilized during training, with $N_{s}=2995$ and $K=5$. During testing, inference is conducted using all candidate ads for each sample, where $N=100,000$ and $K=5$ are employed.
@@ 5.1.3. Baselines
We compare EGA-V1 with the following two representative architectures, which are widely used in industry: 1) MCA. The Multi-stage Cascading Architecture is a common design for online advertising systems. To implement this architecture effectively, we employ four representation methods: SASREC (Kang and McAuley, 2018) for recall, DSSM (Huang et al., 2013) for pre-ranking, DIN (Zhou et al., 2018) for ranking, and GSP (Edelman et al., 2007) for auction. 2) FS-LTR. Full Stage Learning to Rank (Zheng et al., 2024b) is a unified training framework designed for multi-stage recommendation systems. It leverages relabeled data from all stages to train models in MCA, ensuring that the top-ranked items are more likely to pass through subsequent stages and align with user interests. To ensure effectiveness and comparability of FS-LTR, we use the same architecture as MCA, incorporating FS-LTR’s sample re-labeling technique during training to enhance consistency.
@@ 5.1.4. Hyperparameters
For MCA and FS-LTR, we follows the hyperparameters setting in (Liu et al., 2024). For EGA-V1, we tried different hyperparameters using grid search. Due to space limitations, only the most optimal parameters are presented in this paper. The clustering approach $G$ employs an adaptive clustering mechanism, the hidden layers of the MLP are 128 and 32, the learning rate is $10^{-3}$, the batch size is $128$ and the optimizer is Adam. the user behavior sequence length $L=1000$, the dimension size $d=128$, the layer number of GCF and MIF $m=6$, the layer number of RecFormer $m_{e}=3$, the number of interval layers in MIF $m_{k}=2$, the number of attention heads $N_{h}=4$ and the number of adaptive clusters $N_{c}=128$.
@@ 5.2. Offline Experiments
Table 1 summarizes the results on industrial dataset from the offline experiments. Each offline experiment is repeated $5$ times with different random seeds and each result is presented in the form of mean ± standard. The experimental results yield the following observations. On the industrial dataset, EGA-V1 improves over the state-of-the-art baselines in AUC, Recall@50, eCTR, eRPM, and $\Psi$ respectively, demonstrating the superiority of EGA-V1. Specifically, EGA-V1 achieves a Recall@50 of 0.513, representing a significant lift of 20.4% [^5] Lift percentage means the improvement of EGA-V1 over the best baselines over the strongest baseline FS-LTR. The AUC increases to 0.754 (+1.48%), and eCTR improves +8.3%, demonstrating superior user modeling and ranking capabilities. In terms of platform revenue, EGA-V1 achieves an eRPM of 217.1, which is 11.4% higher than FS-LTR. Most notably, the IC metric $\Psi$ is dramatically reduced to 2.3%, compared to 9.1% and 9.4% for FS-LTR and MCA respectively, indicating much stronger incentive compatibility and robustness to auction constraints.
@@ 5.3. Ablation Study
To verify the effectiveness of EGA-V1’s various design considerations, we construct three variants:
[…]
@@ 5.4. Cross Features Analysis
EGA-V1 0.754 0.005
EGA-V1 w/o CF 0.751 (-0.3%) 0.011 (+0.6%)
EGA-V1-${}_{\text{-CA}}$ w/o CF 0.745 (-0.9%) 0.036 (+3.1%)
EGA-V1-${}_{\text{-TA}}$ w/o CF 0.743(-1.1%) 0.047 (+4.2%)
EGA-V1-${}_{\text{-mif}}$ w/o CF 0.737 (-1.7%) 0.048 (+4.4%)
@@ 5.5. Scaling Law
In EGA-V1, each module uses stacked blocks based on cluster-attention as core structure to ensure scalability. Here, we vary the number of blocks across modules to study scaling laws. Besides, we Results in Figure 5 reveal:
@@
2) The figure on the right reveals a clear positive correlation between computational cost and model performance in EGA-V1. The upward-trending linear fit (brown dashed line) indicates that increased computational resources consistently enhance recommendation accuracy, though the marginal gains gradually diminish at higher FLOPs levels. This suggests that while scaling computation remains an effective strategy for performance improvement within the tested range, the model’s architecture or data constraints may eventually limit further gains—mirroring the saturation pattern observed in block stacking. The relationship underscores the importance of balancing computational investment against diminishing returns in practical deployment scenarios.
@@ 5.6. Online Results
To verify EGA-V1’s effectiveness in the real-world, we compare EGA-V1 with the fully deployed MCA in industrial advertising system through online A/B tests. Table 4 presents the results of online A/B testing conducted from November 18 to November 24, 2024 on a low-traffic ad slot of Meituan. Experimental results demonstrate 5.2% improvements in CTR, 13.6% in RPM, and 3.1% in ROI respectively, indicating EGA-V1’s effectiveness in enhancing user experience, platform revenue, and advertiser utility. Notably, the harmonized gains in user engagement and advertiser’s ROI suggest that revenue growth stems not from payment inflation, but from enhanced modeling capabilities.
We also analyzed the online inference computational complexity of EGA-V1 and MCA. EGA-V1 processes hundreds of times more candidate ads than MCA’s ranking model with only a 5 ms average increase (2.2% relatively) in online response time (RT) per request). The complexity is discussed in Appendix A.1.
@@ 6. CONCLUSIONS
In this paper, we propose EGA-V1 that Unifies the entire ad ranking and allocation pipeline as One Model. Our key innovations include: 1) a hybrid feature service that decouples user and ad feature processing to reduce latency while preserving expressiveness; 2) an transformative recommendation framework that leverages an innovative cluster-attention mechanism to efficiently model users’ deep interests and contextual externalities; and 3) a bi-stage training strategy align with user preference and platform objectives. Extensive experiments on public and industrial datasets demonstrate EGA-V1’s superiority over state-of-the-art approaches. In the future, we will further explore scaling laws and different training paradigms to enhance EGA-V1’s capabilities.
@@ A.1. Complexity Analysis
We provide a detailed comparison of computational complexity between MCA and EGA-V1. All floating-point operations (FLOPs) are calculated based on standard Transformer block operations (Vaswani et al., 2017). Given sequence length $L$, hidden dimension $d$, the total FLOPs of a standard Transformer block consist of three components:
@@
EGA-V1 Complexity. The FLOPs of EGA-V1 contains three modules:
@@
Finally, the total Flops of EGA-V1 is calcalated as:
(33) $\text{FLOPs}_{{EGA-V1}}=$ $\text{FLOPs}_{{gcf}}+\text{FLOPs}_{{mif}}+\text{FLOPs}_{{auf}}.$
Both the computational complexities of MCA and EGA-V1 are proportional to the number of candidates($N$). With $N_{r}=\alpha N$, $m_{r}=m$, and $m=2m_{c}$, when $N$ is large, the ratio of proportions can be approximated as follows:
$\frac{\text{FLOPs}_{{EGA-V1}}}{\text{FLOPs}_{{mca}}}\approx\frac{2m_{k}NLd^{2}}{m_{r}N_{r}4L^{2}d}$
In our case, with $\alpha=0.033$, this results in a ratio of $0.97$. The slight difference in proportions leads to a significant FLOPs gap as the number of candidates increases, demonstrating EGA-V1’s capability and efficiency within our design.